The displacement of a particle varies with time (t) as: s = at2 – bt3. The acceleration at any given time (t) of the particle will be equal to


        (a)    2880 m     


(b)    1440 m


        (c)    400 m


        (d)    20 m

Grade:10

1 Answers

Jitender Pal
10 years ago

(b)
s = at2 - bt3
v = ds/dt = 2at - 3bt2
    a=dv/dt  =2a-6bt
    2a-  6bt =0  =>  t=a/3b